# Stability of Jordan Recurrent Neural Network Estimator

### v1

@@ III-A Training and validation
where $1\leq i\leq n$ and $n$ is the number of states, $x^{(t+1)}_{i}$ represents the true value of state $x_{i}$ at time-step $t+1$ and $\hat{x}^{(t+1)}_{i,\phi}$ represents the estimated value of state $x_{i}$ at time-step $t+1$ by the JRN where $\phi$ denotes hyper-parameters of the network. Adam optimization is used for training the network. The optimal learning rate is chosen based on the network’s performance over a range of learning rates. The number of hidden units considered is $50$ and a batch size of $40$ is used for each example in Section IV.
Early stopping with a fixed patience value to decide the number of epochs needed is used. In each epoch, network’s performance on validation data is used as the measure to decide whether to proceed or keep training for another epoch.
III-B Testing and comparison with KF and EKF
We tested our estimators for 3 different systems in Section IV by comparing the results with KF for the linear system and EKF for the nonlinear systems. Standard implementations of KF and EKF are used, see for example [2].
As mentioned earlier, $10\%$ of the data set is used for testing. We compare JRN and EKF graphically using average error at time $t$ over all features and all test sequences
@@ III-B Testing and comparison with KF and EKF
where $m_{test}$ is the number of test sequences, $n$ is the number of states, $x_{i}^{(t)[k]}$ represents the true value of state $x_{i}$ at time-step $t$ for the $k^{th}$ sequence and $\hat{x}_{i}^{(t)[k]}$ represents the estimated value of state $x_{i}$ at time-step $t$ for the $k^{th}$ sequence. We also use normalised mean square error (NMSE) to compare different methods. It is calculated as
$\text{NMSE}=\frac{1}{m_{test}nT}\sum_{k=1}^{m_{test}}\sum_{j=1}^{T}\sum_{i=1}^{n}(x_{i}^{(j)[k]}-\hat{x}_{i}^{(j)[k]})^{2},$ (25)
@@
IV Numerical results
In this section, we illustrate the proposed method with several examples. All three discrete-time systems were obtained by discretizing common ordinary differential equation examples. We use a zero order hold discretization for linear systems. For nonlinear continuous-time state space systems, we use the RK-45 discretization using Python’s $scipy.integrate.solve\_ivp$ function.
A remote Windows server with T4 GPU, courtesy of math faculty computing facility (MFCF), University of Waterloo, was used for all numerical simulations.
(a) Comparison between JRN and EKF Error$(t)$ values for time $0\leq t\leq 50$ seconds for down pendulum.
(b) Comparison between JRN and EKF Error$(t)$ values for time $0\leq t\leq 20$ seconds for reversed Van der Pol oscillator.
Figure 3: The graphs represent a comparison of Error$(t)$ values of Jordan recurrent network-based state estimation and extended Kalman filter-based state estimation for nonlinear dynamical systems. The magnified graph within each graph provides a closer look at the error values from time $19$ to $20$ seconds. The x-axis of both the original and the magnified graphs represents time in seconds, while the y-axis represents error at time $t$. The blue dotted line graph represents the error values for JRN-based estimates, while the orange solid line represents the error values for extended Kalman filter-based estimates.
IV-A Mass-Spring-Damper system
Consider a zero order hold discretization with $\Delta t=0.1s$ of the mass-spring damper system
$\boldsymbol{\dot{x}}(t)=\begin{bmatrix}\dot{x_{1}}(t)\\ \dot{x_{2}}(t)\end{bmatrix}=\begin{bmatrix}x_{2}(t)\\ -\frac{k}{m}x_{1}(t)-\frac{b}{m}x_{2}(t)\end{bmatrix},$ (26)
where $x_{1}(t)$ and $x_{2}(t)$ are the position and velocity respectively of the spring. The parameters are $m=10kg,$ $b=6kg/sec$ and $k=800kg/{sec}^{2}.$ The measurement vector at time-step $t$ is $y^{(t)}=x_{1}^{(t)}+\nu^{(t)}$ where $\nu^{(t)}$ is the measurement noise at time-step $t$.
We consider $100$ sequences from 00 to $100$ seconds each. The mean of the initial condition is chosen randomly from $[-1,1]\times[-1,1].$ Early stopping was implemented with a patience of $5$. We use Adam optimization for training with a learning rate of $10^{-4}.$ There were $4216$ epochs for training.
Fig. 2 shows the error at each time-step averaged all test sequences for JRN and EKF. As can be seen from Fig. 2, the error for JRN converges to 00 as fast as for the KF. Also, the NMSE for JRN is $0.0631$ and for KF is $0.0653$ which is very close. Recall that KF is the optimal estimator for this system. Since the initial condition is not assumed to be known by the estimator, there is a spike in the initial values of the error for the KF and the JRN but the spike in KF error is considerably larger than that of JRN. The magnified graph in Fig. 2 shows that after convergence JRN error is very close to that the KF error.
For this 2-dimensional linear system, the origin is asymptotically stable. We then implement the method in II-A for linear systems. Setting $Q=I$, it is easy to establish that $P=\begin{bmatrix}1.0236&-0.0934\\ -0.0934&69.4428\end{bmatrix}$ With this ISS-Lyapunov function, by Theorem II.6, both the original system and the error system are asymptotically stable. This proves the stability of the NN estimator error.
IV-B Down pendulum
Consider an RK-45 discretization with $\Delta t=0.01s$ of the down pendulum
$\boldsymbol{\dot{x}}(t)=\begin{bmatrix}\dot{x_{1}}(t)\\ \dot{x_{2}}(t)\end{bmatrix}=\begin{bmatrix}x_{2}(t)\\ -\frac{g}{l}\sin({x_{1}(t)})-\frac{b}{m}x_{2}(t)\end{bmatrix},$ (27)
where $x_{1}(t)$ and $x_{2}(t)$ are the position and velocity of the pendulum respectively. The parameters are $g=9.8m/{sec}^{2},$ $m=2kg,$ $b=0.9kg/sec$ and $l=1m.$ The measurement vector at time-step $t$ is $y^{(t)}=x_{1}^{(t)}+\nu^{(t)}$ where $\nu^{(t)}$ is the measurement noise at time-step $t$.
We consider $200$ sequences from 00 to $50$ seconds each. The mean of the initial condition is chosen randomly from $[-2,2]\times[-2,2].$ Early stopping was implemented with a patience of $50$. The validation loss reached early stopping patience at $2984$ epochs. We further used Adam optimization for training with a learning rate of $10^{-4}.$
The error at each time-step, averaged over all features and all test sequences, is shown for JRN and EKF in Fig. 3(a). The error for JRN converges to 00 as fast as that for the EKF. Also, the NMSE for JRN is $0.0627$ while for EKF it is $0.0647$ The magnified graph in Fig. 3(a) shows that after convergence JRN error and EKF error are quite close. The initial spike in JRN error is not as high as the EKF error. The error becomes very small as $t$ goes to $\infty$, despite the presence of noise.
For this nonlinear system, an ISS Lyapunov function was learned using a neural network, as shown in Fig. 4, which is verified by an SMT solver, dReal [27], on a $[-4,4]$ for the error state. In this case, we used $\alpha_{1}(\cdot)=\alpha_{3}(\cdot)=0.01|\cdot|$ and $\alpha_{2}(\cdot)=\gamma(\cdot)=100|\cdot|$.
Figure 4: The learned ISS Lyapunov function for the down pendulum example.
IV-C Reversed Van der Pol oscillator
Consider an RK-45 discretization with $\Delta t=0.1s$ of the reversed Van der Pol oscillator
$\boldsymbol{\dot{x}}(t)=\begin{bmatrix}\dot{x_{1}}(t)\\ \dot{x_{2}}(t)\end{bmatrix}=\begin{bmatrix}-x_{2}(t)\\ x_{1}(t)+(({x_{1}}(t))^{2}-1)x_{2}(t)\end{bmatrix}$ (28)
where $x_{1}(t)$ and $x_{2}(t)$ is the position and velocity of the oscillator. The measurement vector at time-step $t$ is $y^{(t)}=x_{1}^{(t)}+\nu^{(t)}$ where $\nu^{(t)}$ is the measurement noise at time-step $t$.
We consider $200$ sequences from 00 to $20$ seconds each. The mean of the initial condition is chosen randomly from $[-1,1]\times[-1,1].$ The training was done for $35000$ epochs as early stopping patience of $7000$ was not achieved. The validation error continued to decrease. Adam optimization is used for training with a learning rate of $10^{-4}.$ As can be seen from Fig. 3(b), the JRN error converges to zero much faster than the EKF error. Also, the NMSE for JRN is $0.0484$ while for EKF is $0.1025$ This is considerable difference. The magnified graph shows that even after convergence, the JRN error is better than EKF error. This example suggests that when a network is trained for a larger number of epochs, the JRN error can converge much faster than the EKF error.
For this system, in a similar manner to the previous example, we synthesized an ISS-Lyapunov function on $[-2,2]$ for the error states, which provides the asymptotic stability for the error term by Theorem II.6.
V Conclusions and future work
@@
In this paper, we demonstrated that when knowledge of functions $f$ and $h$ in (1) is available or the data of the real systems can be properly sampled, state estimation can be performed by using Jordan recurrent networks. The Jordan recurrent network based estimator helped us achieve results as good as KF and EKF. We also showed the stability of the error dynamics of the neural estimator using an ISS approach, by taking advantage of the structural properties of an unbiased JRN. For linear systems, quadratic Lyapunov equations can be computed analytically for the error dynamics to provide stability. For nonlinear systems, a counter-example guided method was used to learn an ISS Lyapunov function. Furthermore, the effectiveness of the proposed method has been illustrated with three 2-dimensional systems. We will extend our approach to higher dimensional systems in future research. Another potential future research direction is to consider stability for the case where biases are non-zero. Additional further work is to extend the results on the convergence of discrete-time recurrent neural networks as shown in [28] to JRNs.

### v2

@@ IV Implementation
where $1\leq i\leq n$ and $n$ is the number of states, $x^{(t+1)}_{i}$ represents the true value of state $x_{i}$ at time-step $t+1$ and $\hat{x}^{(t+1)}_{i,\phi}$ represents the estimated value of state $x_{i}$ at time-step $t+1$ by the JRN where $\phi$ denotes the weights of the network.
Hyperparameter tuning was performed for all examples individually. Adam optimization is used for training the network. The optimal learning rate is chosen based on the network’s performance on validation dataset over a range of learning rates varying from $10^{-1}$ to $10^{-4}$. A batch size of $40$ and a hidden unit size of $50$ are considered for each example in Section V. Early stopping with a fixed patience value to decide the number of epochs needed is used. In each epoch, network’s performance on validation data is used as the measure to decide whether to proceed or keep training for another epoch. Maximum number of epochs is set to $600$ in case early stopping is not reached. For each lower validation loss value than the previous, the model is saved and after training reloaded to the one corresponding to the least validation loss.
We test our estimators for 3 different systems in Section V. We compare the results with KF for the linear system as it is an optimal filter for systems with additive white Gaussian noise. For nonlinear systems with additive white Gaussian noise, the best filters are considered to be EKF and UKF. Hence, we compare our nonlinear systems results with them. Standard implementations of these 3 filters are used, see for example [1]. For EKF, the equilibrium point around which the systems are linearized is considered to be the origin.
We compare JRN, EKF and UKF graphically using average error at time-step $t$ over all features and all test sequences
@@
where $m_{test}$ is the number of test sequences, $n$ is the number of states, $x_{i}^{(t)[k]}$ represents the true value of state $x_{i}$ at time-step $t$ for the $k^{th}$ sequence and $\hat{x}_{i}^{(t)[k]}$ represents the estimated value of state $x_{i}$ at time-step $t$ for the $k^{th}$ sequence. We also use root mean square error (RMSE) to compare different methods. It is calculated as
$\text{RMSE}=\sqrt{\frac{1}{m_{test}nT}\sum_{k=1}^{m_{test}}\sum_{j=1}^{T}\sum_{i=1}^{n}(x_{i}^{(j)[k]}-\hat{x}_{i}^{(j)[k]})^{2}},$ (23)
@@
Figure 2: Figures 2(a), 2(b) and 2(c) show the training and validation values of JRNs for the mass spring damper system, down pendulum and reversed Van der Pol oscillator respectively. Figures 2(d) 2(e) and 2(f) show average errors at each time-step for JRN, EKF and UKF state estimates for mass spring damper system, down pendulum and reversed Van der Pol oscillator respectively.
Figure 3: Learned ISS Lyapunov function for a down pendulum.
TABLE I: Training hyperparameters
State space system Time-steps Learning rate Patience
Mass spring 300 0.01 10
Down pendulum 300 0.01 25
r. Van der Pol 200 0.001 25
TABLE II: Parameters
State space system Parameters
Mass spring $m=10$ kg, $b=6$ kg/sec, $k=800$ kg/sec2
Down pendulum $m=2$ kg, $b=0.9$ kg/sec, $l=1m.$
V Examples
In this section, we illustrate the proposed method with 3 discrete-time systems obtained as explained in IV. The ODEs considered are mass-spring damper system, down pendulum and reversed Van der Pol oscillator. The measurement vector at time-step $t$ for each of the systems is $y^{(t)}=x_{1}^{(t)}+\nu^{(t)}$ where $x_{1}(t)$ refers to the position and $\nu^{(t)}$ is the measurement noise at time-step $t$. Important parameters for each of these systems are summarised in Tables I and II.A remote Ubuntu server was used for all calculations. The codes used are available on Github at https://github.com/avneetkaur96/JRNs-for-state-estimation.
TABLE III: Root mean square error (RMSE) for E(KF), UKF and JRN
State space Estimator
system E(KF) UKF JRN
Mass spring 0.5132 - 0.5268
Down pendulum 0.1594 0.1693 0.1298
r. Van der Pol 0.1153 0.1158 0.1102
TABLE IV: Best epoch and training time for JRN
State space system Best Epoch Training time(in sec)
Mass spring 135 20
Down pendulum 313 101
r. Van der Pol 96 32
The performance of the JRN and the KF for mass-spring system state estimation is summarized in Figures 2(d) and 2(a). As shown in Figure 2(d), both the JRN and KF achieve rapid error reduction within the first few time steps, with the KF maintaining slightly lower error overall as expected due to its optimal nature. The JRN closely tracks the KF’s performance after the initial steps, indicating effective learning of the system dynamics. The difference of estimation in the first few timesteps can be attributed to the fact that the KF is provided with an estimate of the initial condition while the JRN has no information of the initial condition. The training and validation loss curves for the JRN (Figure 2(a)) demonstrate fast convergence, with loss stabilizing after approximately 40 epochs. The close alignment of training and validation loss throughout training indicates good generalization and minimal overfitting. The RMSE on the test set is $0.5132$ for the KF and $0.5268$ for the JRN (Table III, confirming that the JRN provides comparable estimation accuracy. The model was restored to its state at the best epoch for testing. For this system, the origin is asymptotically stable. We then implement the method for stability of linear systems. Setting $Q=I$, it is easy to establish that $P=\begin{bmatrix}45.504&-3.5737\\ -3.5737&1.4461\end{bmatrix}.$ With this ISS-Lyapunov function, by Theorem III.6, both the original system and the error system are asymptotically stable. This proves the stability of the NN estimator error.
As shown in Figures 2(e) and 2(f), all three approaches exhibit a rapid reduction in error within the first 25 time-steps, stabilizing thereafter.JRN achieves the lowest estimation error over time, outperforming both EKF and UKF for the down pendulum. The errors for both systems and for all methods stabilize after approximately $50$ time-steps, with JRN maintaining a consistently lower error throughout for the down pendulum. Figures 2(b) and 2(c) demonstrate the convergence of the JRN training process. Both training and validation losses decrease steadily, indicating good generalization and absence of overfitting. The JRN achieves the lowest RMSE compared to the EKF and UKF for both the nonlinear systems, indicating improved accuracy (See Table III). These results highlight the effectiveness and efficiency of the JRN in nonlinear state estimation tasks. For both nonlinear systems, an ISS Lyapunov function was learned using a neural network and verified by an SMT solver. For the down pendulum, this function is shown in Fig. 3, verified by an SMT solver, dReal [27], on a $[-2,2]\times[-2,2]$ for the error states. In this case, we used $\alpha_{1}(\cdot)=\alpha_{3}(\cdot)=0.01|\cdot|$ and $\alpha_{2}(\cdot)=\gamma(\cdot)=100|\cdot|$. This provides the asymptotic stability for the error dynamics by Theorem III.6.
VI Conclusions and future work
The Jordan recurrent network based estimator achieved results as good as KF for the linear system and better than EKF and UKF for the chosen nonlinear systems. While the training time is a drawback in case of neural networks, the testing time of JRN-based estimators is much lower than the classical approaches of UKF and EKF. Current work focuses on extending the Jordan architecture to a long short-term neural network [14].
@@ VI Conclusions and future work
The main contribution of this paper is to verify the stability of the error dynamics of the RNN-based estimator using an ISS approach, by taking advantage of the structural properties of an unbiased JRN and recent results in learning Lyapunov functions. For linear systems, quadratic Lyapunov equations can be computed analytically for the error dynamics to provide stability. For nonlinear systems, a counter-example guided method was used to learn an ISS Lyapunov function. The linear approach is used near the equilibrium. The effectiveness of the method has been illustrated with three examples.
Future work will extend this approach to higher-order systems. Addressing the issue of estimation of unstable systems is also of interest. Another potential future research direction is to consider stability for the case where biases are non-zero, and for networks with long short term memory.